# GigaSpeech 2: An Evolving, Large-Scale and Multi-domain ASR Corpus for Low-Resource Languages with Automated Crawling, Transcription and Refinement

### v1

@@ Abstract
The evolution of speech technology has been spurred by the rapid increase in dataset sizes. Traditional speech models generally depend on a large amount of labeled training data, which is scarce for low-resource languages. This paper presents GigaSpeech 2, a large-scale, multi-domain, multilingual speech recognition corpus. It is designed for low-resource languages and does not rely on paired speech and text data. GigaSpeech 2 comprises about 30,000 hours of automatically transcribed speech, including Thai, Indonesian, and Vietnamese, gathered from unlabeled YouTube videos. We also introduce an automated pipeline for data crawling, transcription, and label refinement. Specifically, this pipeline uses Whisper for initial transcription and TorchAudio for forced alignment, combined with multi-dimensional filtering for data quality assurance. A modified Noisy Student Training is developed to further refine flawed pseudo labels iteratively, thus enhancing model performance. Experimental results on our manually transcribed evaluation set and two public test sets from Common Voice and FLEURS confirm our corpus’s high quality and broad applicability. Notably, ASR models trained on GigaSpeech 2 can reduce the word error rate for Thai, Indonesian, and Vietnamese on our challenging and realistic YouTube test set by 25% to 40% compared to the Whisper large-v3 model, with merely 10% model parameters. Furthermore, our ASR models trained on Gigaspeech 2 yield superior performance compared to commercial services. We believe that our newly introduced corpus and pipeline will open a new avenue for low-resource speech recognition and significantly facilitate research in this area.
@@ 1 Introduction
In recent years, the scaling of model parameters and data size has prevailed and proven effective in a range of areas, including language [24, 18], vision [3, 11], as well as speech processing [34, 50, 37]. Consequently, pursuing superior AI models is now closely associated with expanding model size and leveraging larger, high-quality datasets. In the realm of Automatic Speech Recognition (ASR), several large-scale open-source labeled speech datasets [6, 22, 49, 14, 35, 1] have been proposed. However, these extensive datasets are only available for several mainstream languages, such as English and Mandarin, hindering speech recognition development for low-resource languages. Moreover, traditional ASR corpus [1, 10, 5, 12] construction relies heavily on human-labeled speech data, making it time-consuming and a major bottleneck in the fast-paced AI industry. Reducing dependence on vast labeled data is crucial when expanding to new languages and domains [19]. YODAS [29] attempts to address this issue by building multilingual datasets via scraping audio and transcriptions from YouTube. However, neither manual nor automatic subtitles accurately reflect the speech content, resulting in unguaranteed quality.
With this perspective in mind, we propose a new paradigm for constructing large-scale ASR datasets, focusing exclusively on audio content irrespective of the existence or quality of corresponding text pairs. This approach leverages the gigantic amount of unlabeled audio data, thereby bypassing the constraints of scarce paired data. We introduce GigaSpeech 2  [^1] https://huggingface.co/datasets/speechcolab/gigaspeech2 , an evolving, large-scale, multi-domain, multilingual ASR corpus for low-resource Southeast Asian languages. GigaSpeech 2 raw comprises about 30,000 hours of automatically transcribed speech, including Thai, Indonesian, and Vietnamese. GigaSpeech 2 refined consists of 10,000 hours of Thai, 6,000 hours each for Indonesian and Vietnamese. To achieve this, an automated pipeline is developed for data crawling, transcription, and filtering. Furthermore, a modified Noisy Student Training (NST) [47] method is proposed to refine labels from flawed data iteratively. Through comprehensive evaluations, ASR models trained on GigaSpeech 2 refine can reduce the word error rate for Thai, Indonesian, and Vietnamese on our YouTube test set by 25% to 40% compared to the powerful Whisper large-v3 model, with merely 10% model parameters.
@@
  • We release GigaSpeech 2, an evolving, large-scale, multi-domain, and multilingual ASR corpus focusing on low-resource languages. GigaSpeech 2 raw comprises about 30,000 hours of automatically transcribed speech across Thai, Indonesian, and Vietnamese. GigaSpeech 2 refine consists of 10,000 hours of Thai, 6,000 hours each for Indonesian and Vietnamese.
@@
  • We propose a modified NST method to refine flawed pseudo labels iteratively. Our modified NST considers scaling, relabeling, and filtering data within each iteration, significantly improving data quality.
@@ 2 Related Work
Table 1: Comparison of data size between GigaSpeech 2 and other common public multilingual speech datasets on Thai (“th”), Indonesian (“id”), and Vietnamese (“vi”).
Dataset Language Total Duration (h) Domain Speech Type Labeled Label Type
Common Voice [1] th 172.0 Open domain Read Yes Manual
@@
FLEURS [10] th 13.3 Wikipedia Read Yes Manual
@@
VoxLingua107 [44] th 61.0 YouTube Spontaneous No -
@@
CMU Wilderness [4] th 15.6 Religion Read Yes Manual
@@
BABEL [13] vi 87.1 Conversation Spontaneous Yes Manual
VietMed [27] vi 16.0 Medical Spontaneous Yes Manual
Thai Dialect Corpus [41] th 840.0 Open domain Read Yes Manual
TITML-IDN [40] id 14.5 News Read Yes Manual
MEDISCO [36] id 10.0 Medical Read Yes Manual
YODAS manual [29] th 497.1 YouTube Spontaneous Yes Manual
@@
YODAS automatic [29] th 1.9 YouTube Spontaneous Yes Pseudo
@@
Multilingual Low-Resource Speech Datasets Several public multilingual speech datasets have emerged for low-resource languages. BABEL [13], a pioneering dataset, includes conversational telephone data in 17 African and Asian languages. Common Voice [1] offers 19,000 hours of validated recordings in over 100 languages. FLEURS [10] covers 102 languages with 12 hours of supervised data per language. CMU Wilderness [4] provides 20 hours of New Testament data for over 700 languages. VoxLingua107 [44] contains 6,628 hours of unlabeled YouTube data across 107 languages. However, most public multilingual speech datasets focus on high-resource languages, leaving low-resource languages with limited annotated speech data. For example, the available open-source data for Thai, Indonesian, and Vietnamese is scarce, as detailed in Table 1. In contrast, industry-utilized speech models like Whisper [37], MMS [34], Google USM [50], and Universal-1 [38] are trained on massive industrial-grade datasets, the details of which remain undisclosed. To resolve the problem, YODAS [29] attempts to crawl audio from YouTube, but neither manual nor automatic subtitles accurately reflect the speech content, resulting in unguaranteed quality. Moreover, widely used evaluation benchmarks for low-resource languages [1, 10] only consist of read speech, which is relatively clean and mismatched with real-world speech data.
Multilingual Automatic Speech Recognition As the demand for communication between people worldwide grows, many works [37, 50, 34, 28, 30, 43, 8, 33, 42, 23, 9] have shifted attention to multilingual speech recognition. Whisper [37], built on 680,000 hours of web data, supports 99 languages. Google USM [50], trained on YouTube audio, extends to 100+ languages. Massively Multilingual Speech (MMS) [34], trained on religion data, further scales to 1,107 languages.
Noisy Student Training (NST) NST [47, 32, 7, 51] is a self-training technique that leverages unlabeled data to enhance performance. Traditional NST methods start with training a teacher model on high-quality labeled data. Each student model then trains on both noisy-augmented labeled data and pseudo-labeled data generated by its teacher from the unlabeled data.
@@ 3 Dataset Construction
Our proposed automated construction pipeline is illustrated in Fig. 1. Sec. 3.1 covers the stages involved in building GigaSpeech 2 raw and Sec. 3.2 further construct GigaSpeech 2 refined.
@@ 3.1 GigaSpeech 2 raw: Automated Crawling and Transcription
Audio Collection Due to the scarcity of human-labeled data in low-resource languages, our dataset is collected with a focus solely on the audio content, irrespective of the existence or quality of corresponding text pairs. This strategy allows for a broader range of audio data. Given the scarcity and uneven distribution of resources for low-resource languages, we strategically focus on crawling videos from YouTube channels based on two key assumptions. First, prioritizing popular channels ensures consistent domain characteristics and audio quality. Second, different channels have no speaker overlap, simplifying the subsequent data partitioning. The data collection process starts by manually defining categories of interest. The selected topics include Agriculture, Art, Business, Climate, Culture, Economics, Education, Entertainment, Health, History, Literature, Music, Politics, Relationships, Shopping, Society, Sport, Technology, and Travel. Alongside multiple topics, various content formats are also considered, including Audiobook, Commentary, Lecture, Monologue, Movie, News, Talk, and Vlog. This broad selection ensures the comprehensiveness of the dataset across multiple domains for research and analysis. Once the list of YouTube channels is prepared, we use yt-dlp  [^2] https://github.com/yt-dlp/yt-dlp toolkit to download all audio files in WebM format. These files are then converted to WAV format with a single channel and resampled at a 16 kHz sampling rate.
Creating TRAIN/DEV/TEST Splits To ensure no speaker overlap between the splits, we manually verify that there is no speaker overlap between different channels and partition the data by allocating different YouTube channels to each subset. The dataset is divided into three distinct subsets: training (TRAIN), development (DEV), and TETS. The DEV and TEST sets each contain 10 hours and are manually transcribed by professionals, while the remainder is allocated to the TRAIN set. Table 1 shows the amount of data across these three languages. Detailed analysis of GigaSpeech2 is illustrated in Appendix B.
Transcription with Whisper Whisper large-v3 model  [^3] https://huggingface.co/openai/whisper-large-v3 from OpenAI is employed to transcribe audio files automatically. For each audio recording, a 30-second segment is selected from the middle to perform language detection using Whisper. Only recordings that match the target languages are transcribed.
Forced Alignment with TorchAudio Although Whisper can generate timestamps, inspection reveals they are not precise enough. We resort to the model  [^4] https://dl.fbaipublicfiles.com/mms/torchaudio/ctc_alignment_mling_uroman/model.pt from TorchAudio [20] for forced alignment, which provides reliable alignment for noisy transcriptions, supports efficient processing on GPUs, and handles longer sequences more effectively [34].
@@
Multi-dimensional Filtering A series of heuristic filtering rules across text and audio modalities are implemented to exclude relatively poor-quality samples.
  • Charset Filtering: Segments are retained if they only contain characters permitted by the charset of the respective language.
  • Language Confidence Filtering: The language identification (LID) model  [^5] https://dl.fbaipublicfiles.com/fasttext/supervised-models/lid.176.bin from fastText [21] is used to filter based on the estimated language confidence score, retaining only segments with confidence scores above a predetermined threshold. This method effectively eliminates meaningless and repetitive segments. Note that language identification based on audio has already been performed before transcription.
  • Audio Duration Filtering: Segments are filtered based on duration, with only those retained within the predetermined minimum and maximum duration thresholds.
  • Balancing: We carefully control the duplication of transcripts caused by channel-specific content while preserving natural linguistic patterns.
@@ 3.2 GigaSpeech 2 refined: Iterative Label Refinement
Some samples remain low quality due to inaccuracies in Whisper transcriptions and imprecise forced alignment boundaries. To address this, we develop a modified NST method. As illustrated in the bottom right corner of Fig. 1, it starts by training a teacher model on a subset of flawed pseudo labels and iteratively expands the training set, generates new pseudo labels, and filters them. A student model, equal to or larger than the teacher, is trained on refined pseudo labels and assigned as the new teacher. SpecAugment [31], Bypass [48], and feature mask [48] introduce noise during each NST step. Bypass, a type of stochastic depth, learns channel-wise scalar weights to combine the module input and output. Feature mask performs dropout in the hidden dimension of the feedforward and convolution layer but shares over the time dimension. This deliberate noising allows the student model to learn consistency with the teacher model, which remains unaffected by noise when generating pseudo labels [47]. This iterative process gradually enhances data quality. Detailed algorithm steps are shown in Appendix A Algo. 1.

### v2

@@ Abstract
The evolution of speech technology has been spurred by the rapid increase in dataset sizes. Traditional speech models generally depend on a large amount of labeled training data, which is scarce for low-resource languages. This paper presents GigaSpeech 2, a large-scale, multi-domain, multilingual speech recognition corpus. It is designed for low-resource languages and does not rely on paired speech and text data. GigaSpeech 2 comprises about 30,000 hours of automatically transcribed speech, including Thai, Indonesian, and Vietnamese, gathered from unlabeled YouTube videos. We also introduce an automated pipeline for data crawling, transcription, and label refinement. Specifically, this pipeline involves Whisper for initial transcription, MMS for forced alignment, and multi-dimensional filtering for data quality assurance. A modified Noisy Student Training is developed to further refine flawed pseudo labels iteratively, thereby enhancing model performance. Experimental results on our manually transcribed evaluation set and two public test sets from Common Voice and FLEURS confirm our corpus’s high quality and broad applicability. Notably, ASR models trained on GigaSpeech 2 can reduce the word error rate for Thai, Indonesian, and Vietnamese on our challenging and realistic YouTube test set by 25% to 40% compared to Whisper large-v3, with merely 10% model parameters. Furthermore, our ASR models trained on GigaSpeech 2 yield superior performance compared to commercial services. We hope that our newly introduced corpus and pipeline will open a new avenue for low-resource speech recognition and significantly facilitate research in this area.
GigaSpeech 2: An Evolving, Large-Scale and Multi-domain ASR Corpus for Low-Resource Languages with Automated Crawling, Transcription and Refinement
Yifan Yang1, Zheshu Song1, Jianheng Zhuo1, Mingyu Cui3 Jinpeng Li4, Bo Yang2, Yexing Du2,6, Ziyang Ma1, Xunying Liu3, Ziyuan Wang7 Ke Li8, Shuai Fan9, Kai Yu1,9, Wei-Qiang Zhang4,11, Guoguo Chen10,11, Xie Chen1,5,11thanks: Corresponding author. Email: chenxie95@sjtu.edu.cn 1X-LANCE Lab, SCS, MoE Key Lab of Artificial Intelligence, Shanghai Jiao Tong University 2Pengcheng Laboratory 3The Chinese University of Hong Kong 4Dept EE, Tsinghua University 5SII 6Harbin Institute of Technology 7Birch AI 8Dataocean AI 9AISpeech Ltd 10Seasalt AI Inc 11SpeechColab
@@ 1 Introduction
In recent years, the scaling of model parameters and data size has prevailed and proven effective in a range of areas, including language Kaplan et al. (2020); Hoffmann et al. (2022), vision Betker et al. (2023); Dehghani et al. (2023), as well as speech processing Pratap et al. (2024); Zhang et al. (2023); Radford et al. (2023). Consequently, pursuing superior AI models is now closely associated with expanding model size and leveraging larger, high-quality datasets. In the realm of Automatic Speech Recognition (ASR), several large-scale open-source labeled speech datasets Chen et al. (2021); Kang et al. (2024); Zhang et al. (2022); Galvez et al. (2021); Pratap et al. (2020b); Ardila et al. (2020) have been proposed. However, these extensive datasets are only available for several mainstream languages, such as English and Mandarin, hindering speech recognition development for low-resource languages. Moreover, traditional ASR corpus Ardila et al. (2020); Conneau et al. (2023); Bu et al. (2017); Du et al. (2018) construction relies heavily on human-labeled speech data, making it time-consuming and a major bottleneck in the fast-paced AI industry. Reducing dependence on vast labeled data is crucial when expanding to new languages and domains. YODAS Li et al. (2023) attempts to address this issue by building multilingual datasets via scraping audio and transcriptions from YouTube. However, neither manual nor automatic subtitles accurately reflect the speech content, resulting in unguaranteed quality.
With this perspective in mind, we propose a new paradigm for constructing large-scale ASR datasets, focusing solely on audio content irrespective of the existence or quality of corresponding text pairs. This approach leverages the gigantic amount of unlabeled audio data, bypassing the constraints of scarce paired data. We introduce GigaSpeech 2, an evolving [^1] The term “evolving” continues the naming convention used by GigaSpeech., large-scale, multi-domain, multilingual ASR corpus for low-resource Southeast Asian languages. GigaSpeech 2 raw comprises about 30,000 hours of automatically transcribed speech, across Thai, Indonesian, and Vietnamese. GigaSpeech 2 refined consists of 10,000 hours of Thai, 6,000 hours each for Indonesian and Vietnamese. To achieve this, an automated pipeline is developed for data crawling, transcription, and filtering. Furthermore, a modified Noisy Student Training (NST) Xie et al. (2020) method is proposed to refine labels from flawed data iteratively. Through comprehensive evaluations, ASR models trained on GigaSpeech 2 refined can reduce the word error rate for Thai, Indonesian, and Vietnamese on our YouTube test set by 25% to 40% compared to the powerful Whisper large-v3 model, with merely 10% model parameters.
@@
  • We release GigaSpeech 2 with two versions: GigaSpeech 2 raw comprises about 30,000 hours of automatically transcribed speech across Thai, Indonesian, and Vietnamese. GigaSpeech 2 refined consists of 10,000 hours of Thai, 6,000 hours each for Indonesian and Vietnamese.
@@
  • We propose a modified NST method to iteratively refine flawed pseudo labels. Our modified NST performs scaling, relabeling, and filtering data within each iteration, significantly improving final data quality.
@@ 2 Related Work
Table 1: Comparison of data size between GigaSpeech 2 and other common public multilingual speech datasets on Thai (th), Indonesian (id), and Vietnamese (vi).
Dataset Language # Hours (h) Domain Speech Type Labeled Label Type
Common Voice Ardila et al. (2020) th 172.0 Open domain Read Yes Manual
@@
FLEURS Conneau et al. (2023) th 13.3 Wikipedia Read Yes Manual
@@
VoxLingua107 Valk and Alumäe (2021) th 61.0 YouTube Spontaneous No -
@@
CMU Wilderness Black (2019) th 15.6 Religion Read Yes Manual
@@
BABEL Gales et al. (2014) vi 87.1 Conversation Spontaneous Yes Manual
VietMed Le-Duc (2024) vi 16.0 Medical Spontaneous Yes Manual
Thai Dialect Corpus Suwanbandit et al. (2023) th 840.0 Open domain Read Yes Manual
TITML-IDN Shinoda and Furui (2011) id 14.5 News Read Yes Manual
MEDISCO Qorib and Adriani (2018) id 10.0 Medical Read Yes Manual
YODAS manual Li et al. (2023) th 497.1 YouTube Spontaneous Yes Manual
@@
YODAS automatic Li et al. (2023) th 1.9 YouTube Spontaneous Yes Pseudo
@@
Multilingual Low-Resource Speech Datasets Several publicly available multilingual speech datasets have emerged for low-resource languages. BABEL Gales et al. (2014), a pioneering dataset, includes conversational telephone data in 17 African and Asian languages. Common Voice Ardila et al. (2020) offers 19,000 hours of validated recordings in over 100 languages. FLEURS Conneau et al. (2023) covers 102 languages with 12 hours of supervised data per language. CMU Wilderness Black (2019) provides 20 hours of New Testament data for over 700 languages. VoxLingua107 Valk and Alumäe (2021) contains 6,628 hours of unlabeled YouTube data across 107 languages. However, most public multilingual speech datasets focus on high-resource languages, leaving low-resource languages with limited annotated speech data. As detailed in Table 1, the available open-source data for Thai, Indonesian, and Vietnamese is scarce. In contrast, industry-utilized speech models like Whisper Radford et al. (2023), MMS Pratap et al. (2024), Google USM Zhang et al. (2023), and Universal-1 Ramirez et al. (2024) are trained on massive industrial-grade datasets, the details of which remain undisclosed. To resolve the problem, YODAS Li et al. (2023) attempts to crawl audio from YouTube, but neither manual nor automatic subtitles accurately reflect the speech content, resulting in unguaranteed quality. Moreover, widely used evaluation benchmarks for low-resource languages Ardila et al. (2020); Conneau et al. (2023) only consist of read speech, which is relatively clean and mismatched with real-world speech data.
Multilingual Automatic Speech Recognition As the demand for communication between people worldwide grows, many works Radford et al. (2023); Zhang et al. (2023); Pratap et al. (2024); Li et al. (2021); Lugosch et al. (2022); Toshniwal et al. (2018); Cho et al. (2018); Pratap et al. (2020a); Tjandra et al. (2023); Kannan et al. (2019); Conneau et al. (2021) have shifted attention to multilingual speech recognition. Whisper Radford et al. (2023), built on 680,000 hours of web data, supports 99 languages. Google USM Zhang et al. (2023), trained on YouTube audio, extends to 100+ languages. Massively Multilingual Speech (MMS) Pratap et al. (2024), trained on religion data, further scales to 1,107 languages.
Noisy Student Training (NST) NST Xie et al. (2020); Park et al. (2020); Xu et al. (2020); Zhang et al. (2020); Likhomanenko et al. (2021); Mehmood et al. (2022); Chen et al. (2023) is a self-training technique that leverages unlabeled data to enhance performance. Traditional NST methods start with training a teacher model on high-quality labeled data. Each student model then trains on both noisy-augmented labeled data and pseudo-labeled data generated by its teacher from the unlabeled data. One study Chen et al. (2023) has explored using Character Error Rate (CER), calculated between pseudo-labeled data generated with and without language model, to perform data selection, suggesting a positive correlation between the CERs of different pseudo labels and their ground truth.
@@ 3 Dataset Construction
Our proposed automated construction pipeline is illustrated in Fig. 1. Sec. 3.1 covers the stages involved in building GigaSpeech 2 raw and Sec.  3.2 further construct GigaSpeech 2 refined.
@@ 3.1 GigaSpeech 2 raw: Automated Crawling and Transcription
Audio Collection Due to the scarcity of human-labeled data in low-resource languages, our dataset is collected with a focus solely on the audio content, irrespective of the existence or quality of corresponding text pairs. This strategy allows for leveraging a broader range of audio data. Given the scarcity and uneven distribution of resources for low-resource languages, we strategically crawl videos from YouTube channels based on two key considerations. First, prioritizing mainstream and popular channels helps ensure consistent domain characteristics and higher audio quality. Such content is widely viewed, and its creators are generally more mindful of ethical and legal considerations prior to publishing. Second, channels with huge differences in topics and content formats are less likely to have speaker overlap, which simplifies subsequent data partitioning. The data collection process starts by manually defining categories of interest. The selected topics include Agriculture, Art, Business, Climate, Culture, Economics, Education, Entertainment, Health, History, Literature, Music, Politics, Relationships, Shopping, Society, Sport, Technology, and Travel. Alongside multiple topics, various content formats are also considered, including Audiobook, Commentary, Lecture, Monologue, Movie, News, Talk, and Vlog. This broad selection ensures the comprehensiveness of the dataset across multiple domains for research and analysis. Once the list of YouTube channels is prepared, we use yt-dlp [^2] https://github.com/yt-dlp/yt-dlp toolkit to download all audio files in WebM format. These files are then converted to WAV format with a single channel and resampled at a 16 kHz sampling rate.
Creating TRAIN/DEV/TEST Splits To ensure no speaker overlap between the splits, we manually verify no speaker overlap between different channels and partition the data by allocating different YouTube channels to each subset. The dataset is divided into three distinct subsets: TRAIN, DEV, and TEST. The DEV and TEST sets each contain 10 hours and are manually transcribed by professionals, while the remainder is allocated to the TRAIN set. Table 1 shows the amount of data across these three languages. Detailed analysis of GigaSpeech 2 is illustrated in Appendix A.
Transcription with Whisper Whisper large-v3 model [^3] https://huggingface.co/openai/whisper-large-v3 from OpenAI is used to transcribe audio files automatically. For each audio recording, a 30-second segment is selected from the middle to perform language detection by Whisper. Only audios that match the target languages are transcribed.
Forced Alignment with TorchAudio Although Whisper can generate timestamps, inspection reveals they are not precise enough. We resort to the model [^4] https://dl.fbaipublicfiles.com/mms/torchaudio/ctc_alignment_mling_uroman/model.pt from TorchAudio Hwang et al. (2023) for forced alignment, which provides reliable alignment for noisy transcriptions, supports efficient processing on GPUs, and handles longer sequences more effectively (Pratap et al., 2024).
@@
Multi-dimensional Filtering A series of heuristic filtering rules across text and audio modalities are implemented to exclude relatively poor-quality samples. 1) Charset Filtering: Segments are retained if they only contain characters permitted by the charset of the respective language. 2) Language Confidence Filtering: The language identification (LID) model [^5] https://dl.fbaipublicfiles.com/fasttext/supervised-models/lid.176.bin from fastText (Joulin et al., 2016) is used to filter based on the estimated language confidence score, retaining only segments with confidence scores above a predetermined threshold. This method effectively eliminates meaningless and repetitive segments. Note that language identification based on audio has already been performed before transcription. 3) Audio Duration Filtering: Segments are filtered based on duration, with only those retained within the predetermined minimum and maximum duration thresholds. 4) Balancing: We carefully control the duplication of transcripts caused by channel-specific content while preserving natural linguistic patterns.
@@ 3.2 GigaSpeech 2 refined: Iterative Label Refinement
Some samples remain low quality due to inaccuracies in Whisper transcriptions and imprecise forced alignment boundaries. To address this, we develop a modified NST method. As illustrated in the bottom right corner of Fig. 1, it begins by training a teacher model on a subset of flawed pseudo labels, iteratively expanding the training set, generating new pseudo labels, and filtering them. A student model, equal to or larger than the teacher, is trained on these refined pseudo labels and assigned as the new teacher. Unlike previous NST approaches that heavily rely on unchanged supervised data combined with additional unsupervised data, our method eliminates the need for any supervised data. Instead, we treat the flawed pseudo labels generated by Whisper as supervised data, refining all labels iteratively based on the Character Error Rate (CER) between those produced by Whisper and the teacher model. SpecAugment (Park et al., 2019), Bypass (Yao et al., 2024), and feature mask (Yao et al., 2024) introduce noise during each NST step. Bypass, a type of stochastic depth, learns channel-wise scalar weights to combine the module input and output. Feature mask performs dropout in the hidden dimension of the feedforward and convolution layer but shares across the time dimension. This deliberate noising enables the student model to learn consistency with the teacher model, which remains unaffected by noise when generating pseudo labels (Xie et al., 2020). This iterative process progressively enhances data quality. Algo. 1 illustrates the workflow of our proposed iterative label refinement.
Output: Refined-label set $\mathcal{R}$
$\mathcal{R}\leftarrow\mathcal{P}_{1}$;
Train teacher model $\mathcal{M}_{1}$ on $\mathcal{R}$ with noise;
for $i\leftarrow 1$ to $n$ do
       $\mathcal{R}\leftarrow\varnothing$;
       if $i$ == $1$  then
             // Filter $\mathcal{P}_{i}$ by teacher model $\mathcal{M}_{i}$ with CER $\leq\tau$
      else
             for $j\leftarrow 1$ to $i$ do
                   $\mathcal{R}\leftarrow\mathcal{R}\cup\mathcal{R}_{tmp}$;
             end for
       end if
end for
return $\mathcal{R}$;
Algorithm 1 Iterative Label Refinement